OpenStudy (anonymous):

the altitude of a triangle is increasing at a rate of 1 cm/min while the area in increasing at a rate of 2 cm^2/min. At what rate is the base of the triangle changing when the altitude is 10 cm and the area is 100cm^2?

hero (hero):

You got this

hero (hero):

What's the formula we use here?

OpenStudy (anonymous):

is there a shorcut? i feel like i just have to double something then square it.

hero (hero):

1. Draw diagram, include given info 2. Apply proper formula 3. Take derivative of formula 4. Input values into formula 5. Isolate what you need to find

OpenStudy (anonymous):



hero (hero):

I skipped a step which is to find other necessary values

OpenStudy (amistre64):

h': 1 cm/min A': 2 cm^2/min. b'?: h=10 cm, A=100cm^2 A = bh/2 A' = b'h/2 + bh'/2 b' = (A' - bh'/2)(2/h)

OpenStudy (anonymous):

given: dA/dt= 1 dAREA/dt=2 A=10 AREA = 100

OpenStudy (amistre64):

define b as 2A/h

hero (hero):

Amistre, go to bed, lol

OpenStudy (anonymous):

im gona try and work it out then ill shoot you guys an answer. give me a sec

OpenStudy (amistre64):

b' = (A' - 2Ah'/2h)(2/h) b' = (A' - Ah'/h)(2/h) b' = (2 - 100(1)/10) (2/10) b' = (2 - 10) (1/5) b' = -8/5 maybe?

OpenStudy (amistre64):

..... i thought this was bed :)

hero (hero):

I believe you

OpenStudy (anonymous):

why is b, 2A/h

OpenStudy (amistre64):

becasue A = bh/2; Area = base*height/2 ... solveing for base we get; base = 2*Area/height